a bathtub takes 10 minutes to fill with water and 12 minutes to drain. If the drain stays open unnoticed, how long will it take to fill the tub with water?
Answer by bluemockingjay7(41)

rate x time = work
Fill rate = 1/10
Time = x
Work = 1/10 * x = x/10
Drain rate = 1/12
Time = x
Work = 1/12 * x = x/12
x/10 - x/12 = 1
LCM = 60
60(x/10 - x/12 = 1)
6x - 5x = 60
x = 60
It will take 60 minutes or 1 hour to fill the tub with water.